# Phase error rate estimation with basis-efficiency mismatch for decoy-state BB84

### v1

@@ V.2.1 Testing Rounds after basis-independent Filter
V.2.2 Key Generation Rounds after basis-independent Filter
@@
Since the measurements in the key generation rounds leading to $\tilde{e}^{\text{key}}_{ZX}$ are not identical to the one in the testing rounds which leads to $\tilde{e}^{\text{obs}}_{XX}$, one cannot directly use Serfling (Lemma 2) to relate the two, as we did in Section IV.3. Therefore, we introduce another set of virtual measurements (which take place in the Key* ($X\rightarrow X$) box in Fig. 2), corresponding to $X$ basis second and third step measurements. Thus we obtain another error rate $\tilde{e}^{\text{key}}_{XX}$. This is the error rate corresponding to the case where these $\tilde{n}_{K}$ rounds are measured using $\Big{\{}G^{\text{con,F}}_{(XX),(\neq)},$ $G^{\text{con,F}}_{(XX),(=)},\mathrm{I}-{F}_{(X),(\text{con})}\Big{\}}$ (the same measurement that testing rounds are subject to).
V.3 Cost of removing the basis-independent loss assumption
@@
Consider the POVM elements $G^{\text{con,F}}_{(ZX),(\neq)}$ and $G^{\text{con,F}}_{(XX),(\neq)}$ defined via Eq. 27, which combine the second and third step measurements. In Section IV they were exactly equal. We define $\delta_{1}$ to quantify the closeness of these POVM elements as
@@
Consider the second step measurements, where outcomes corresponding to POVM element $\mathrm{I}-{F}_{(Z),(\text{con})}$ are discarded. In Section IV, there was no need of the second step filtering measurement, which is equivalent to having ${F}_{(Z),(\text{con})}=\mathrm{I}$. We define $\delta_{2}$ to quantify the amount of deviation from this case as
@@ V.4 Sampling
    $e^{\text{obs}}_{X}$ is the fraction of the $n_{X}$ testing rounds that resulted in the $G^{\text{con,F}}_{(XX),(\neq)}$ outcome. We have access to $e^{\text{obs}}_{X}$ in the protocol, since it is something we actually observe.
@@
  • $\bm{\tilde{e}^{\text{obs}}_{XX}}\leftrightarrow\bm{\tilde{e}^{\text{key}}_{XX}}$ : These error rates correspond to measurement outcomes using the same POVM, but with random assignment to testing vs key generation. Thus we can apply Lemma 2 (Serfling) in exactly the same manner as in Section IV.3, conditioned on the event $\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}$. In doing so, we obtain
@@
  • $\bm{\tilde{e}^{\text{key}}_{XX}}\leftrightarrow\bm{\tilde{e}^{\text{key}}_{ZX}}$: We utilize the definition of $\delta_{1}$ stated in Section V.3. Since the POVM elements generating $\bm{\tilde{e}^{\text{key}}_{ZX}}$ ($G^{\text{con,F}}_{(ZX),(\neq)}$) and $\bm{\tilde{e}^{\text{key}}_{XX}}$ ($G^{\text{con,F}}_{(XX),(\neq)}$) are close, we expect the bounds obtained on $\bm{\tilde{e}^{\text{key}}_{ZX}}$ and $\bm{\tilde{e}^{\text{key}}_{XX}}$ to also be close. This is made precise in the following lemma proved in Section B.2.
@@
    $\Pr(\bm{\tilde{e}^{\text{key}}_{ZX}}\geq f+\delta_{1}+\gamma^{\varepsilon_{\text{AT-b}}}_{\text{bin}}(\tilde{n}_{K},\delta_{1}))_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}\leq\Pr(\bm{\tilde{e}^{\text{key}}_{XX}}\geq f)_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}+\varepsilon_{\text{AT-b}}^{2}.$ (38)
@@
    Recall that $\bm{n_{K}}$ is obtained by discarding rounds from $\bm{\tilde{n}_{K}}$ based on $\{{F}_{(Z),(\text{con})},\mathrm{I}-{F}_{(Z),(\text{con})}\}$ measurements. We will essentially show that very few rounds are discard in this step, using Eq. 29. To do so, we prove the following Lemma in Appendix B.
@@
Thus we have relationships Eqs. 30, 33, 38 and 41 between all the error rates that hold with high probability. These can all be combined using straightforward but cumbersome algebra (see Appendix C), to obtain
@@
Thus we now have a phase error estimation bound that is valid even in the presence of basis-efficiency mismatch. We summarize the results of this section in the following theorem. Note that the results of Section IV are a special case ($\delta_{1},\delta_{2}=0$) of the following theorem.
@@
V.5 Relation to other work on basis-efficiency mismatch
@@
Ref. [12] proposed the first security proof of QKD in the presence of Eve controlled basis-efficiency mismatch, in the asymptotic regime for IID collective attacks. It required the assumption that a qubit is received on Bob’s side, for which it required the use of detector-decoy methods. Furthermore it argued that the basis-efficiency mismatch can be removed entirely (for qubits received by Bob) if one randomly swaps the 00 and the $1$ detector. Note that this trick only works for the qubit subspace and does not hold for higher photon numbers. Ref. [11] extended this work without having to assume qubit detections, but still performs an IID asymptotic analysis. In Ref. [39], the numerical framework for key rate computations was proposed and used to compute IID asymptotic key rates for perfectly characterized and fixed basis-efficiency mismatch. In Ref. [13] the numerical framework [39] was used to compute IID asymptotic key rates for a toy model where Eve was allowed to induce basis efficiency mismatch via the use of two spatio-temporal modes. Recently, Ref. [14] used the numerical framework to estimate the phase error rate in the presence of basis efficiency mismatch, but again assumed IID collective attacks in the asymptotic regime, and single qubits received on Bob’s side. The assumption of single qubits received by Bob was later removed in the follow up work [15]. In both [14, 15], Eve is not allowed to control the efficiency mismatch.
@@
VI Application to Decoy-state BB84
@@
First, we rigorously prove security for variable-length protocols (note that Ref. [28] actually implicitly implements a variable-length protocol). In fact [28] considers a protocol with iterative sifting  [^6] This has been formulated in a variety of ways in the literature. In general, we use this phrase for protocols that have a “loop” phase, where some actions are taken repeatedly until certain conditions are met. where the total number of rounds in the protocol is not fixed a priori, and depends on observations made during the protocol. However a justification for this step is not provided. This is important because certain kinds of iterative sifting can lead to subtle issues in the security analysis [38, 46]. We do not fix this problem directly in this work. Instead, we consider a protocol with a fixed total number of signals sent, which avoids this problem. Second, we make rigorous certain technical steps in [28] (regarding entropic calculations on states conditioned on events), which we point out where applicable. Third, our phase error estimates do not require the assumption of basis-independent loss unlike that of [28]. We also avoid a particular Taylor series approximation used by [28] ([47, Eq. 22]), and therefore our phase error estimation procedure yields a true bound without any approximations. Finally we also clarify certain aspects of the decoy analysis undertaken in [28]. In particular, we careful differentiate between random variables and an observed value of the random variable, and also properly condition on relevant events in our presentation.
@@
Recently, a more accessible version of the security proof in Ref. [28] was written in Ref. [48]. We note that while Ref. [48] is certainly quite pedagogical, it does not address the issues described above. We stress that while we clarify and make rigorous certain steps in [28], our main contribution in this work is the variable-length security proof of decoy-state BB84 in the presence of basis-efficiency mismatch.
We start by first specifying the decoy-state BB84 protocol we study in Section VI.1. We will then explain the required bounds on the phase error rate in Section VI.2. We explain decoy analysis in Section VI.3, and prove the security of our variable-length protocol in Section VI.5. Some proofs are delegated to Appendix D.
VI.1 Protocol specification
@@
The remaining steps of the protocol are the same as in Section II. In particular, based on the observations $n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}$, Alice and Bob implement one-way error-correction using $\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$ bits of communication, followed by error-verification, and privacy amplification to produce a key of $l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$ bits. (This requirement of one-way communication is not fundamental, and more complicated error-correction protocols can be accommodated in a straightforward manner (see Remark 4). Additionally note that our protocol generates key from all intensities, instead of having a single “signal” intensity for key generation.
VI.2 Required and actual phase error estimation bound
@@
where $\mathcal{B}^{\text{decoy}}_{\text{min}-m}$ and $\mathcal{B}^{\text{decoy}}_{\text{max}-m}$ are functions that compute bounds on the $m$-photon components of the input statistics. Note that we use $\bm{n_{X_{\neq},\mu_{\vec{k}}}}=(\bm{n_{X_{\neq},\mu_{1}}}\times\bm{e^{\text{obs}}_{X,\mu_{1}}},\dots,\bm{n_{X_{\neq},\mu_{3}}}\times\bm{e^{\text{obs}}_{X,\mu_{3}}})$ to denote the number of rounds resulting both Alice and Bob using the $X$ basis and obtaining an error, for each intensity (and we will assume implicit conversion between these two notations). The $9$ on the RHS comes from the fact that we implement decoy analysis on $3$ different events and we have $3$ intensities. We will prove Eq. 49 in Section VI.3.
@@
In this second step, we will use $\bm{e^{\text{obs}}_{X,1}},\bm{n_{X,{1}}},\bm{n_{K,{1}}}$ to bound the single photon phase error rate $\bm{e^{\text{key}}_{X,{1}}}$. Notice this is exactly what we showed Sections IV and V. In particular, with $\mathcal{B}_{\delta_{1},\delta_{2}}$ directly obtained from Eq. 44, we have
@@
However, note that we do not directly observe $\bm{e^{\text{obs}}_{X,1}},\bm{n_{X,{1}}},\bm{n_{K,{1}}}$ in the decoy-state protocol (unlike Section V). Thus we would like to replace these values with the bounds computed from our decoy analysis (Eq. 49). This is straightforward to do, since $\mathcal{B}_{\delta_{1},\delta_{2}}$ is an increasing function of $\bm{e^{\text{obs}}_{X,1}}$, and decreasing function of $\bm{n_{X,{1}}},\bm{n_{K,{1}}}$. This can be done formally by a straightforward application of the union bound for probabilities ($\Pr(\Omega_{1}\lor\Omega_{2})\leq\Pr(\Omega_{1})+\Pr(\Omega_{2})$) applied to Eqs. 49 and 50. Doing so allows us to conclude that the probability of any of the bounds in Eqs. 49 and 50 failing is smaller than $9\varepsilon_{\text{AT-d}}^{2}+\varepsilon_{\text{AT-s}}^{2}$. Then we use the fact that if none of the bounds inside the probabilities in Eqs. 49 and 50 fail, then this implies that the bounds inside the probability in Eq. 51 below must hold. Formally, we obtain
@@ VI.3 Decoy Analysis
The application of Hoeffdings inequality here is subtle, and is made rigorous in Lemmas 12 and 13 in Appendix D. In particular, we do not claim that the photon numbers of the above pulses (pulses that led to outcome $O$) are sampled independently, or that intensities of the above pulses are chosen independently of one another (they are not). Rather, Lemmas 12 and 13 rely on exploiting the fact that conditioned on any fixed sequence of photon numbers of the pulses, the intensities are chosen independently of one another. Basically, we first need to condition on a fixed sequence of photon numbers, apply Hoeffdings on the intensity choices, and then get rid of the conditioning. One cannot apply Hoeffdings directly.
@@
VI.4 Bounds on zero and one photon statistics
For any event $O\in\{X,X_{\neq},K\}$, the relevant bounds on the zero-photon and single-photon components can be obtained by algebraic manipulation of the expressions in $\mathcal{S}_{\text{constraints}}$. In general, any method for bounding the relevant zero-photon and single-photon components using $\mathcal{S}_{\text{constraints}}$ suffices. In this work, we follow exactly the steps taken by Ref. [28, Appendix A] to obtain these bounds. Thus, we only write the final expressions here. We define

### v2

@@ IV.2.1 Testing Rounds after basis-independent Filter
IV.2.2 Key Generation Rounds after basis-independent Filter
@@
Since the measurements in the key generation rounds leading to $\tilde{e}^{\text{key}}_{ZX}$ are not identical to the one in the testing rounds which leads to $\tilde{e}^{\text{obs}}_{XX}$, one cannot directly use Serfling (Lemma 2) to relate the two, as we did in Section III.3. Therefore, we introduce another set of virtual measurements (which take place in the Key* ($X\rightarrow X$) box in Fig. 3), corresponding to $X$ basis second and third step measurements. Thus we obtain another error rate $\tilde{e}^{\text{key}}_{XX}$. This is the error rate corresponding to the case where these $\tilde{n}_{K}$ rounds are measured using $\Big{\{}G^{\text{con,F}}_{(XX),(\neq)},$ $G^{\text{con,F}}_{(XX),(=)},\mathrm{I}-{F}_{(X),(\text{con})}\Big{\}}$ (the same measurement that testing rounds are subject to).
IV.3 Cost of removing the basis-independent loss assumption
@@
Consider the POVM elements $G^{\text{con,F}}_{(ZX),(\neq)}$ and $G^{\text{con,F}}_{(XX),(\neq)}$ defined via Eq. 17, which combine the second and third step measurements. In Section III they were exactly equal. We define $\delta_{1}$ to quantify the closeness of these POVM elements as
@@
Consider the second step measurements, where outcomes corresponding to POVM element $\mathrm{I}-{F}_{(Z),(\text{con})}$ are discarded. In Section III, there was no need of the second step filtering measurement, which is equivalent to having ${F}_{(Z),(\text{con})}=\mathrm{I}$. We define $\delta_{2}$ to quantify the amount of deviation from this case as
@@ IV.4 Sampling
    $e^{\text{obs}}_{X}$ is the fraction of the $n_{X}$ testing rounds that resulted in the $G^{\text{con,F}}_{(XX),(\neq)}$ outcome. We have access to $e^{\text{obs}}_{X}$ in the protocol, since it is something we actually observe. (Recall that this is the only error rate we observe).
@@
  • $\bm{\tilde{e}^{\text{obs}}_{XX}}\leftrightarrow\bm{\tilde{e}^{\text{key}}_{XX}}$ : These error rates correspond to measurement outcomes using the same POVM, but with random assignment to testing vs key generation. Thus we can apply Lemma 2 (Serfling) in exactly the same manner as in Section III.3, conditioned on the event $\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}$. In doing so, we obtain
@@
  • $\bm{\tilde{e}^{\text{key}}_{XX}}\leftrightarrow\bm{\tilde{e}^{\text{key}}_{ZX}}$: We utilize the definition of $\delta_{1}$ stated in Section IV.3. Since the POVM elements generating $\bm{\tilde{e}^{\text{key}}_{ZX}}$ ($G^{\text{con,F}}_{(ZX),(\neq)}$) and $\bm{\tilde{e}^{\text{key}}_{XX}}$ ($G^{\text{con,F}}_{(XX),(\neq)}$) are close, we expect the bounds obtained on $\bm{\tilde{e}^{\text{key}}_{ZX}}$ and $\bm{\tilde{e}^{\text{key}}_{XX}}$ to also be close. This is made precise in the following lemma proved in Section C.2.
@@
    $\Pr(\bm{\tilde{e}^{\text{key}}_{ZX}}\geq e+\delta_{1}+\gamma^{\varepsilon_{\text{AT-b}}}_{\text{bin}}(\tilde{n}_{K},\delta_{1}))_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}\leq\Pr(\bm{\tilde{e}^{\text{key}}_{XX}}\geq e)_{|\Omega_{(\tilde{n}_{X},\tilde{n}_{K})}}+\varepsilon_{\text{AT-b}}^{2}.$ (28)
@@
    Recall that $\bm{n_{K}}$ is obtained by discarding rounds from $\bm{\tilde{n}_{K}}$ based on $\{{F}_{(Z),(\text{con})},\mathrm{I}-{F}_{(Z),(\text{con})}\}$ measurements. We will essentially show that very few rounds are discard in this step, using Eq. 19. To do so, we prove the following Lemma in Appendix C.
@@
Thus we have relationships Eqs. 20, 23, 28 and 31 between all the error rates that hold with high probability. These can all be combined using straightforward but cumbersome algebra (see Appendix D), to obtain
@@
Thus we now have a phase error estimation bound that is valid even in the presence of basis-efficiency mismatch. We summarize the results of this section in the following theorem. Note that the results of Section III are a special case ($\delta_{1},\delta_{2}=0$) of the following theorem.
@@
Proof sketch. The rigorous proof follows from the analysis already seen in Section IV. Essentially, we consider several error rates corresponding to various measurement choices, as described in Fig. 3, and use Lemmas 2, 3 and 4 to relate the various error rates together.
@@
IV.5 Relation to other work on basis-efficiency mismatch
@@
Ref. [12] proposed the first security proof of QKD in the presence of Eve controlled basis-efficiency mismatch, in the asymptotic regime for IID collective attacks. It required the assumption that a qubit is received on Bob’s side, for which it required the use of detector-decoy methods. Furthermore it argued that the basis-efficiency mismatch can be removed entirely (for qubits received by Bob) if one randomly swaps the 00 and the $1$ detector. Note that this trick only works for the qubit subspace and does not hold for higher photon numbers (see Section VI.3). Ref. [11] extended this work without having to assume qubit detections, but still performs an IID asymptotic analysis. In Ref. [38], the numerical framework for key rate computations was proposed and used to compute IID asymptotic key rates for perfectly characterized and fixed basis-efficiency mismatch. In Ref. [13] the numerical framework [38] was used to compute IID asymptotic key rates for a toy model where Eve was allowed to induce basis efficiency mismatch via the use of two spatio-temporal modes. Recently, Ref. [14] improved upon [12] by obtaining tighter estimates on the phase error rate in the presence of basis efficiency mismatch, but again assumed IID collective attacks in the asymptotic regime, and single qubits received on Bob’s side. The assumption of single qubits received by Bob was later removed in the follow up work [15]. In both [14, 15], Eve is not allowed to control the efficiency mismatch. Finally, MDI-QKD [10] addresses all detector side-channels in the finite-size regime against coherent attacks, but requires a drastically different protocol implementation compared to standard QKD.
@@
V Application to Decoy-state BB84
@@
First, we rigorously prove the security for the variable-length decoy-state BB84 protocol in the entropic uncertainty relations framework (note that Ref. [30] actually implicitly implements a variable-length protocol). In fact [30] considers a protocol with iterative sifting  [^5] This has been formulated in a variety of ways in the literature. In general, we use this phrase for protocols that have a “loop” phase, where some actions are taken repeatedly until certain conditions are met. where the total number of rounds in the protocol is not fixed a priori, and depends on observations made during the protocol, which are announced in a round-by-round manner. However a justification for this step is not provided. This is important because certain kinds of iterative sifting can lead to subtle issues in the security analysis (see Ref. [45] for some issues and Ref. [46] for solutions). We do not fix this problem directly in this work. Instead, we consider a protocol with a fixed total number of signals sent, which avoids this problem. Second, we make rigorous certain technical steps in [30] (regarding entropic calculations on states conditioned on events), which we point out where applicable (see Remark 19). Third, our phase error estimates do not require the assumption of basis-independent loss unlike that of [30]. We also avoid a particular Taylor series approximation used by [30] ([47, Eq. 22]), and therefore our phase error estimation procedure yields a true bound without any approximations. Finally we also clarify certain aspects of the decoy analysis undertaken in [30]. In particular, we careful differentiate between random variables and an observed value of the random variable, and also properly condition on relevant events in our presentation. We stress that while we clarify and make rigorous certain steps in [30], our main contribution in this work is the variable-length security proof of decoy-state BB84 in the presence of detector imperfections.
@@
Recently, a more accessible version of the security proof in Ref. [30] was written in Ref. [48]. We note that while Ref. [48] is certainly quite pedagogical, it does not address all the issues described above.
We start by first specifying the decoy-state BB84 protocol we study in Section V.1. We will then explain the required bounds on the phase error rate in Section V.2. We explain decoy analysis in Section V.3, and state the security of our variable-length protocol in Section V.5. Some proofs are delegated to Appendix E.
V.1 Protocol specification
@@
The remaining steps of the protocol are the same as in Section II. In particular, based on the observations $n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z}$, Alice and Bob implement one-way error-correction using $\lambda_{\text{EC}}(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$ bits of communication, followed by error-verification, and privacy amplification to produce a key of $l(n_{X,\mu_{\vec{k}}},n_{K,\mu_{\vec{k}}},e^{\text{obs}}_{X,\mu_{\vec{k}}},e^{\text{obs}}_{Z})$ bits. This requirement of one-way communication is not fundamental, and more complicated error-correction protocols can be accommodated in a straightforward manner (see Remark 4). Additionally note that our protocol generates key from all intensities, instead of having a single “signal” intensity for key generation.
V.2 Required and actual phase error estimation bound
@@
where $\mathcal{B}^{\text{decoy}}_{\text{min}-m}$ and $\mathcal{B}^{\text{decoy}}_{\text{max}-m}$ are functions that compute bounds on the $m$-photon components of the input statistics. Note that we use $\bm{n_{X_{\neq},\mu_{\vec{k}}}}=(\bm{n_{X,\mu_{1}}}\times\bm{e^{\text{obs}}_{X,\mu_{1}}},\dots,\bm{n_{X,\mu_{3}}}\times\bm{e^{\text{obs}}_{X,\mu_{3}}})$ to denote the number of rounds resulting both Alice and Bob using the $X$ basis and obtaining an error, for each intensity (and we will assume implicit conversion between these two notations). The $9$ on the RHS comes from the fact that we implement decoy analysis on $3$ different events and we have $3$ intensities. We will prove Eq. 39 in Section V.3.
@@
In this second step, we will use $\bm{e^{\text{obs}}_{X,1}},\bm{n_{X,{1}}},\bm{n_{K,{1}}}$ to bound the single photon phase error rate $\bm{e^{\text{key}}_{X,{1}}}$. Notice this is exactly what we showed Sections III and IV. In particular, with $\mathcal{B}_{\delta_{1},\delta_{2}}$ directly obtained from Eq. 34, we have
@@
However, note that we do not directly observe $\bm{e^{\text{obs}}_{X,1}},\bm{n_{X,{1}}},\bm{n_{K,{1}}}$ in the decoy-state protocol (unlike Section IV). Thus we would like to replace these values with the bounds computed from our decoy analysis (Eq. 39). This is straightforward to do, since $\mathcal{B}_{\delta_{1},\delta_{2}}$ is an increasing function of $\bm{e^{\text{obs}}_{X,1}}$, and decreasing function of $\bm{n_{X,{1}}},\bm{n_{K,{1}}}$. This can be done formally by a straightforward application of the union bound for probabilities ($\Pr(\Omega_{1}\lor\Omega_{2})\leq\Pr(\Omega_{1})+\Pr(\Omega_{2})$) applied to Eqs. 39 and 40. Doing so allows us to conclude that the probability of any of the bounds in Eqs. 39 and 40 failing is smaller than $9\varepsilon_{\text{AT-d}}^{2}+\varepsilon_{\text{AT-s}}^{2}$. Then we use the fact that if none of the bounds inside the probabilities in Eqs. 39 and 40 fail, then this implies that the bounds inside the probability in Eq. 41 below must hold. Formally, we obtain
@@ V.3 Decoy Analysis
The application of Hoeffdings inequality here is subtle, and is made rigorous in Lemmas 12 and 13 in Appendix E (see also Ref. [43]). Note that in general, the photon numbers of every pulse in the protocol are chosen independently, since Alice chooses intensity independently for each pulse. However, here we are interested in photon numbers corresponding to rounds that led to a specific outcome $O$. Since we postselect pulses based on the outcome, we can no longer claim that the photon numbers of these pulses (pulses that led to outcome $O$) are sampled independently, or that intensities of these pulses are chosen independently. This is because they now depend on Eve’s attack. Rather, Lemmas 12 and 13 rely on exploiting the fact that conditioned on any fixed sequence of photon numbers of the pulses, the intensities are chosen independently of one another. One can therefore apply Hoeffdings inequality. Then, since the resulting statements holds for any fixed sequence of photon numbers, the conditioning on this event can be removed.
@@
V.4 Bounds on zero and one photon statistics
For any event $O\in\{X,X_{\neq},K\}$, the relevant bounds on the zero-photon and single-photon components can be obtained by algebraic manipulation of the expressions in $\mathcal{S}_{\text{constraints}}$. In general, any method for bounding the relevant zero-photon and single-photon components using $\mathcal{S}_{\text{constraints}}$ suffices. In this work, we follow exactly the steps taken by Ref. [30, Appendix A] to obtain these bounds. Thus, we only write the final expressions here. We define